The average speed of a car over a certain distance is calculated by dividing the
total distance travelled by the total time taken. Two towns A and B are 10 km
apart, A motorist travels from A to B at an average speed of 20 km per hour,
and travels from B to A at an average 40 km per hour.
2.3.1 Determine the average speed to the motorist for the two trips
combined. (i.e. from A to B and the back to A) (6)
2.3.2 Determine the speed would the motorist have to average on the return trip from
B to A to yield an overall speed of 30km per hour? Assume the motorist
has averaged 20km per hour from A to B) (4)
2.4 AutoZone sells a regular car battery for R 700 and heavy-duty model for
R1 100. If it sold 40 batteries yesterday for a total of R34 000.
Indicate how many of each of the type of and indicate the rand the rand value sales of
each type.

Answer by mananth(16946)

Two towns A and B are 10 km apart, A motorist travels from A to B at an average speed of 20 km per hour,and travels from B to A at an average 40 km per hour.
2.3.1 Determine the average speed to the motorist for the two trips
combined. (i.e. from A to B and the back to A) (6)
2.3.2 Determine the speed would the motorist have to average on the return trip from
B to A to yield an overall speed of 30km per hour? Assume the motorist
has averaged 20km per hour from A to B) (4)

Two towns A and B are 10 km apart
speed = 20km/h
Time taken = d/t = 10/20 = 1/2 hour
B to A at an average 40 km per hour
Time taken = 10/40 = 1/4 hour
Total time = 1/2 +1/4 = 3/4 hour
Total distance = 10+10 =20 km
Average speed = 20/(3/4) = 80/3 =26.66 km/h
Let the return speed be x km/h
time taken = 10/x
Total time taken will be (1/2)+(10/x)
distance = 20 km
Total average speed required = 30 km/h
20/((1/2 +10/x ) =30
20/((x+20)/2x)= 30
20 * 2x/(x+20) = 30
40x = 30x +600
10x = 600
x= 60 km/h

x Number of regular batteries sold
y Number of special batteries sold


1 x + y = 40


700 x + 1100 y = 34000 .............2
Eliminate y
multiply (1)by -1100
Multiply (2) by 1
-1100 x -1100 y = -44000
700 x 1100 y = 34000
Add the two equations
-400.00 x = -10000
/ -400.00
x = 25
plug value of x in (1)
1.00 x + y = 40
25.00 + y = 40
y = 15
y = 15
Ans x = 25
y = 15
0 25.00 Number of regular batteries sold
0 15.00 Number of special batteries sold
Rest you can do